OpenStudy (anonymous):

do you use PEMDAS for this problem ? 3+6-5+1

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

no you just solve it in the order the numbers appear, so 3+6=9 and then 9-5=4 and then 4+1=5

OpenStudy (anonymous):

how so?

OpenStudy (anonymous):

you will get the same answer.

OpenStudy (anonymous):

3+6=9 9+1=10 10-5=5

OpenStudy (anonymous):

it doesnt really matter the order for addition or subtraction so you can put it any any order you would like, sometimes it is just easier to add or subtract as you go in the order it appears in the question, but if you like to use order of operations that works too!

OpenStudy (amistre64):

the order of subtraction matters ... 5-3 is not the same as 3-5

OpenStudy (amistre64):

but as far as add or subtract first; there only rule is to work it left to right unless you have groupings to define it in a different manner

OpenStudy (anonymous):

thats what i was trying to say

OpenStudy (amistre64):

i know, but some people get the idea that the AS is a rule for addition over subtraction; which is just not the case

OpenStudy (anonymous):

just put it into ur calculator

OpenStudy (anonymous):

you cant always be reliable upon a calculator to solve your problems

OpenStudy (amistre64):

3+6=9 9+1=10 10-5=5 is not the "order" that it is presented in; you in fact changed the equation to: 3+6+(-5)+1, and commuted the 1 and -5 as: 3+6+1+(-5)

OpenStudy (anonymous):

i was using PEMDAS

OpenStudy (amistre64):

not really, since PEMDAS doesnt have a preference of addition over subtraction as far as ORDER of operations is concerned. The rule is to work it left to right. What you did was employ an algebraic device. Spose we did it this way: 3+6-5+1 lets add first (3+6)-(5+1) 9-6 , and now subtract 9-6 = 3

OpenStudy (amistre64):

order of operations is an ORDER in which you work out the operations defined in the mathical sentence ....

OpenStudy (anonymous):

i never thought of it that way

OpenStudy (anonymous):

Good point